OpenStudy (anonymous):

calculate the arc length of the indicated portion of the curve r(t) r(t)=11t^4i+2t^4j+10t^4k 1

OpenStudy (dumbcow):

\[Arc = \int\limits_{1}^{3}\sqrt{1+r'(t)^{2}}dt\] not sure what to do with the i,j,and k

OpenStudy (anonymous):

oh ok thanks for the guidence can you help with more

OpenStudy (dumbcow):

sure

OpenStudy (anonymous):

so i believe the equation to use here is L (arc length) = \[\int\limits_{1}^{3} [r'(t)]dt\]

OpenStudy (anonymous):

you want to find the derivatives of each i, j, and k... so (44t^3)i + (8t^3)k + (40t^3)k after that you evaluate each function... sqrt((dx)^2 + (dy)^2 + (dz)^2)... this essentially squares the r'(t) vector components and puts it under the sqrt.

OpenStudy (anonymous):

after substituting and simplifying you should come out with 60t^3 *assuming i did all the algrebra right!* you then integrate it which becomes 15t^4... finally you apply your limits so 15(3^4)-15(1^4) = 1200.